# VST++: Efficient and Stronger Visual Saliency Transformer

### v1

@@ 1 Introduction
Salient object detection (SOD) aims at identifying and segmenting the most prominent objects in images. It can be utilized as a pre-processing technique to promote various computer vision tasks, such as weakly-supervised semantic segmentation [1], person re-identification [2], and video object segmentation [3, 4, 5]. Apart from RGB images, depth data captured by depth cameras provides valuable spatial structure information. RGB-D SOD capitalizes on these two data types and has garnered increasing attention in recent years.
@@ 2.1 CNN-Based SOD
CNN-based methods have become a prevailing trend in the SOD field, delivering impressive performances owing to the powerful feature learning ability of CNNs. RGB SOD methods can be broadly classified into four groups: attention-based, multi-level fusion-based, recurrent-model-based, and multi-task-based approaches. Multi-level fusion-based approaches [11, 41, 12, 52, 13] often utilized UNet [8] and HED-style [53] networks as a foundation to aggregate both high-level and low-level features. Besides, some works [41, 14] adopted spatial pyramid pooling and [52, 13] applied the dilated convolution [54] for gathering multi-scale features. Attention-based methods [55, 21, 13, 22] and [20] effectively selected discriminative features by resorting to spatial, channel, or pixel-wise contextual attention. Recurrent-model-based works [40, 56, 57, 58, 59] focused on refining the details of saliency map step by step. Multi-task-based methods tried to introduce other tasks, such as fixation prediction [15], image caption [16], and edge detection [60, 17, 18, 48, 19], to enhance SOD performance. For a more comprehensive literature review, please refer to [61].
For RGB-D SOD, an effective fusion of RGB and depth features is critical. For simplicity, early works [26, 62, 27] utilized basic operations (e.g., concatenation, summation, or multiplication) to fuse information from two modalities. Recent studies [63, 64, 55, 23] leveraged attention mechanism such as spatial or channel attention generated from depth cues to improve the RGB features. Additionally, dynamic convolution [29], graph neural networks [14], and knowledge distillation [30] were also used to perform multi-modal feature fusion. As a counterpart, cross-attention was invoked in [42, 65, 22] to propagate long-range cross-modality cues. Please refer to [66] for a more comprehensive literature review.
Another multi-modal SOD task, i.e. RGB-T SOD, also aims to explore complementary cues for accurate salient object detection. Various traditional graph-based approaches have been proposed to obtain low-level hand-crafted features [67, 68, 69]. Recently, some works have used CNNs to extract image features and proposed diverse fusion methods, including multi-level feature fusion [70, 71], spatial-wise attention mechanism [72], and channel attention mechanism module [73].
@@ 2.2 Transformers in Computer Vision
Figure 1: Overall architecture of our proposed VST++ model for both RGB and RGB-D SOD. The encoder generates multi-level tokens from the input image patch sequence. Next, the convertor projects the patch tokens to the decoder space and performs cross-modal information fusion for RGB-D SOD. Finally, a multi-task transformer decoder simultaneously performs saliency detection and boundary segmentation via the proposed task-related tokens and the patch-task-attention mechanism. An RT2T transformation is designed to progressively upsample the patch tokens. In each decoder layer, we design the Select-Integrate Attention (SIA) to select fine-grained foreground segments and aggregate coarse-grained background information for low-cost attention computation. Depth Position Embedding (DPE) or sinusoidal position encoding (PE) is added to the query and key of the SIA for RGB-D SOD and RGB SOD, respectively. Additionally, we calculate the token-supervised prediction losses at every decoder level. The dotted line represents components exclusively designed for RGB-D SOD.
@@ 3.3.1 Token Upsampling and Multi-level Token Fusion
Besides, drawing inspiration from the demonstrated achievements of multi-level feature fusion in existing SOD methods [11, 12, 52, 13, 14], we make use of low-level tokens with larger lengths from the encoder to provide precise local structural details. For both RGB and RGB-D SOD, we exclusively use low-level tokens derived from the RGB transformer encoder. Specifically, we gradually fuse $\bm{T}_{2}$ and $\bm{T}_{1}$ with the upsampled patch tokens through concatenation and linear projection, formulated as:
@@ 3.3.3 Select-Integrate Attention
$\bm{t}^{g}_{i}=\dfrac{{(1-\bm{M}_{i})}{\bm{T}^{\mathcal{D}}_{i}}}{\begin{matrix}\sum_{k=1}^{l_{i}}(1-{\bm{M}_{i}})\end{matrix}}.$ (8)
@@ 3.4 Loss Function
where $\mathcal{L}_{s}$ and $\mathcal{L}_{b}$ represent the saliency prediction loss and the boundary prediction loss in total, respectively. $\widehat{\bm{y}_{s_{i}}}$ and $\widehat{\bm{y}_{b_{i}}}$ denote the ground truth for saliency and boundary tasks, respectively. The $\mathcal{L}_{den}^{i}$ and $\mathcal{L}_{sup}^{i}$ are implemented as BCE losses.
The ultimate loss function is formulated as their summation:
@@ 4.3.1 Effectiveness of Token-supervised Prediction Loss
Summary DUTS[90] ECSSD[91] HKU-IS[92] PASCAL-S[93] DUT-O[95] SOD[96] Metric MACs Params $S_{m}$ maxF $E_{\xi}^{\text{max}}$ MAE $S_{m}$ maxF $E_{\xi}^{\text{max}}$ MAE $S_{m}$ maxF $E_{\xi}^{\text{max}}$ MAE $S_{m}$ maxF $E_{\xi}^{\text{max}}$ MAE $S_{m}$ maxF $E_{\xi}^{\text{max}}$ MAE $S_{m}$ maxF $E_{\xi}^{\text{max}}$ MAE CNN-based PiCANet[20] 54.05 47.22 0.863 0.840 0.915 0.040 0.916 0.929 0.953 0.035 0.905 0.913 0.951 0.031 0.846 0.824 0.882 0.071 0.826 0.767 0.865 0.054 0.813 0.824 0.871 0.073 AFNet[109] 21.66 35.95 0.867 0.838 0.910 0.045 0.914 0.924 0.947 0.042 0.905 0.910 0.949 0.036 0.849 0.824 0.877 0.076 0.826 0.759 0.861 0.057 0.811 0.819 0.867 0.085 TSPOANet[110] - - 0.860 0.828 0.907 0.049 0.907 0.919 0.942 0.047 0.902 0.909 0.950 0.039 0.841 0.817 0.871 0.082 0.818 0.750 0.858 0.062 0.802 0.809 0.852 0.904 EGNet-R[17] 157.21 111.64 0.887 0.866 0.926 0.039 0.925 0.936 0.955 0.037 0.918 0.923 0.956 0.031 0.852 0.825 0.874 0.080 0.841 0.778 0.878 0.053 0.824 0.831 0.875 0.080 ITSD-R[47] 15.96 26.47 0.885 0.867 0.929 0.041 0.925 0.939 0.959 0.035 0.917 0.926 0.960 0.031 0.861 0.839 0.889 0.771 0.840 0.792 0.880 0.061 0.835 0.849 0.889 0.075 MINet-R[12] 87.11 162.38 0.884 0.864 0.926 0.037 0.925 0.938 0.957 0.034 0.919 0.926 0.960 0.029 0.856 0.831 0.883 0.071 0.883 0.769 0.869 0.056 0.830 0.835 0.878 0.074 LDF-R[18] 15.51 25.15 0.892 0.877 0.930 0.034 0.925 0.938 0.954 0.034 0.920 0.929 0.958 0.028 0.861 0.839 0.888 0.067 0.839 0.782 0.870 0.052 0.831 0.841 0.878 0.071 CSF-R2[106] 18.96 36.53 0.890 0.869 0.929 0.037 0.931 0.942 0.960 0.033 - - - - 0.863 0.839 0.885 0.073 0.838 0.775 0.869 0.055 0.826 0.832 0.883 0.079 GateNet-R[52] 162.22 128.63 0.891 0.874 0.932 0.038 0.924 0.935 0.955 0.038 0.921 0.926 0.959 0.031 0.863 0.836 0.886 0.071 0.840 0.782 0.878 0.055 0.827 0.835 0.877 0.079 Transformer-based VST[44] 41.36 44.48 0.896 0.877 0.939 0.037 0.932 0.944 0.964 0.034 0.928 0.937 0.968 0.030 0.873 0.850 0.900 0.067 0.850 0.800 0.888 0.058 0.854 0.866 0.902 0.065 EBMGSOD[84] 53.38 118.96 0.909 0.900 0.949 0.029 0.941 0.954 0.972 0.024 0.930 0.943 0.971 0.023 0.877 0.856 0.899 0.061 0.858 0.817 0.900 0.051 0.866 0.881 0.915 0.056 ICON-B[83] 52.59 94.30 0.917 0.911 0.960 0.024 0.941 0.954 0.971 0.024 0.936 0.947 0.974 0.022 0.885 0.860 0.903 0.058 0.869 0.830 0.906 0.043 0.862 0.878 0.909 0.059 VST-t++ 30.07 46.05 0.897 0.881 0.939 0.035 0.933 0.946 0.966 0.030 0.928 0.938 0.969 0.026 0.869 0.849 0.899 0.066 0.854 0.812 0.894 0.056 0.855 0.870 0.905 0.062 VST-T++ 28.44 53.6 0.901 0.887 0.943 0.033 0.937 0.949 0.968 0.029 0.930 0.939 0.968 0.026 0.878 0.855 0.901 0.063 0.853 0.804 0.892 0.053 0.853 0.866 0.899 0.065 VST-S++ 32.73 74.9 0.909 0.897 0.947 0.029 0.939 0.951 0.969 0.027 0.932 0.941 0.969 0.025 0.880 0.859 0.901 0.062 0.859 0.813 0.890 0.050 0.859 0.866 0.898 0.059 VST-B++ 69.30 112.23 0.923 0.915 0.957 0.024 0.950 0.962 0.977 0.021 0.941 0.951 0.976 0.021 0.889 0.867 0.907 0.056 0.871 0.827 0.900 0.044 0.872 0.886 0.910 0.056
@@ 4.4 Comparison with State-of-the-Art Methods
For RGB SOD, we compare our VST++ with nine state-of-the-art CNN-based RGB SOD methods including PiCANet [20], AFNet [109], TSPOANet [110], EGNet-R [17], ITSD-R [47], MINet-R [12], LDF-R [18], CSF-R2 [106], and GateNet-R [52], as well as three state-of-the-art transformer-based RGB SOD methods: EBMGSOD [84], ICON [83], and our previous VST [44]. Table III shows the comparison results.
For RGB-D SOD, we adopt 14 state-of-the-art CNN-based RGB-D SOD methods, i.e., $S^{2}$ [42], PGAR [59], DANet [111], cmMS [112], ATSA [24], CMW [23], Cas-Gnn [14], HDFNet [29], CoNet [19], BBS-Net [13], JL-DCF-R [27], SPNet [113], CMINet [114], DCF [115] , as well as four state-of-the-art transformer-based RGB-D SOD models, namely SwinNet [85], HRTransNet [86], EBMGSOD [84], and our previous VST [44], for comparison. Table IV and Table V report the comparison results.
[…]

### v2

@@ 1 Introduction
Salient object detection (SOD) aims at identifying and segmenting the most prominent objects in images. It can be utilized as a pre-processing technique to promote various computer vision tasks, such as weakly-supervised learning [1, 2], person re-identification [3], and video object segmentation [4, 5, 6]. Apart from RGB images, depth data captured by depth cameras provides valuable spatial structure information. RGB-D SOD capitalizes on these two data types and has garnered increasing attention in recent years.
@@ 2.1 CNN-Based SOD
CNN-based methods have become a prevailing trend not only in the segmentation field [53], delivering impressive performances owing to the powerful feature learning ability of CNNs. RGB SOD methods can be broadly classified into four groups: attention-based, multi-level fusion-based, recurrent-model-based, and multi-task-based approaches. Multi-level fusion-based approaches [12, 42, 13, 54, 14] often utilized UNet [9] and HED-style [55] networks as a foundation to aggregate both high-level and low-level features. Besides, some works [42, 15] adopted spatial pyramid pooling and [54, 14] applied the dilated convolution [56] for gathering multi-scale features. Attention-based methods [57, 22, 14, 23] and [21] effectively selected discriminative features by resorting to spatial, channel, or pixel-wise contextual attention. Recurrent-model-based works [41, 58, 59, 60, 61] focused on refining the details of saliency map step by step. Multi-task-based methods tried to introduce other tasks, such as fixation prediction [16], image caption [17], and edge detection [62, 18, 19, 49, 20], to enhance SOD performance. For a more comprehensive literature review, please refer to [32]. For RGB-D SOD, an effective fusion of RGB and depth features is critical. For simplicity, early works [27, 63, 28] utilized basic operations (e.g., concatenation, summation, or multiplication) to fuse information from two modalities. Recent studies [64, 65, 57, 24] leveraged attention mechanism such as spatial or channel attention generated from depth cues to improve the RGB features. Additionally, dynamic convolution [30], graph neural networks [15], and knowledge distillation [31] were also used to perform multi-modal feature fusion. As a counterpart, cross-attention was invoked in [43, 66, 23] to propagate long-range cross-modality cues. Please refer to [33] for a more comprehensive literature review.
Another multi-modal SOD task, i.e. RGB-T SOD, also aims to explore complementary cues for accurate salient object detection. Various traditional graph-based approaches have been proposed to obtain low-level hand-crafted features [67, 68, 69]. Recently, some works have used CNNs to extract image features and proposed diverse fusion methods, including multi-level feature fusion [70, 71], spatial-wise attention mechanism [72], and channel attention mechanism module [73]. Moreover, CNN-based architectures have been introduced in the context of light field SOD [74, 75], deep unsupervised SOD [76], and Part-Object Relational Saliency [77].
@@ 2.2 Transformers in Computer Vision
Figure 1: Overall architecture of our proposed VST++ model for both RGB and RGB-D SOD. The encoder generates multi-level tokens from the input image patch sequence. Next, the convertor projects the patch tokens to the decoder space and performs cross-modal information fusion for RGB-D SOD. Finally, a multi-task transformer decoder simultaneously performs saliency detection and boundary segmentation via the proposed task-related tokens and the patch-task-attention mechanism. An RT2T transformation is designed to progressively upsample the patch tokens. In 1/8 and 1/4 decoder levels, we design Select-Integrate Attention (SIA) to select fine-grained foreground segments and aggregate coarse-grained background information for low-cost attention computation. We maintain the self-attention mechanism for the 1/16 decoder level, given the absence of a mask from the previous stage. Depth Position Embedding (DPE) or sinusoidal position encoding (PE) is added to the query and key of the SIA for RGB-D SOD and RGB SOD, respectively. Additionally, we calculate the token-supervised prediction losses at every decoder level. The dotted line represents components exclusively designed for RGB-D SOD.
@@ 2.3 Transformer-Based SOD
2.4 Generic Dense Segmentation Architecture
Recently, several generic dense prediction frameworks have emerged for a range of segmentation tasks. For instance, K-Net [92] introduced distinct dynamic kernels tailored for class-specific segmentation. Within DETR-based models, SOLQ [93] employed object-specific queries to predict class, location, and mask. CMT-DeepLab [94] treated object queries as cluster centers, facilitating the grouping of pixels for segmentation. In Panoptic SegFormer [95], query responsibilities were segregated, directing them to the location decoder and mask decoder. Mask2Former [96], building on MaskFormer [97], integrated masked attention to achieve a universal segmentation design. Expanding upon Mask2Former, FastInst [98] and MP-former [99] were introduced, featuring innovatively designed instance activation-guided queries and a mask-piloted training approach.
While we acknowledge the superiority of generic dense prediction models, it’s important to highlight a substantial gap when applying them to SOD tasks owing to the lack of class information. Unlike the approaches mentioned above, our VST++ designed modules specifically for SOD tasks. For example, our multi-task decoder places emphasis on boundary detection, taking into account the characteristics of the SOD task. The introduction of SIA is geared towards diminishing the computational overhead associated with self-attention, by leveraging the nature of SOD. Additionally, we put forward the DPE method, specifically tailored to encode depth position for RGB-D SOD. These designs helps us achieve superior performance on the SOD task compared with existing general models.
@@ 3.3.1 Token Upsampling and Multi-level Token Fusion
Besides, drawing inspiration from the demonstrated achievements of multi-level feature fusion in existing SOD methods [12, 13, 54, 14, 15], we make use of low-level tokens with larger lengths from the encoder to provide precise local structural details. For both RGB and RGB-D SOD, we exclusively use low-level tokens derived from the RGB transformer encoder. Specifically, we gradually fuse $\bm{T}_{i}$ with the upsampled patch tokens $\bm{T}^{\mathcal{D}}_{i+1}$through concatenation and linear projection, formulated as:
@@ 3.3.3 Select-Integrate Attention
$\bm{t}^{g}_{i}=\dfrac{{(1-\bm{M}_{i})}{\bm{T}^{\mathcal{D}}_{i}}}{\begin{matrix}\sum_{k=1}^{l_{i}}(1-{\bm{M}_{i}[:,k]})\end{matrix}}.$ (8)
@@ 3.4 Loss Function
where $\mathcal{L}_{s}$ and $\mathcal{L}_{b}$ represent the saliency prediction loss and the boundary prediction loss in total, respectively. $\widehat{\bm{y}_{s_{i}}}$ and $\widehat{\bm{y}_{b_{i}}}$ denote the ground truth for saliency and boundary tasks, respectively. The $\mathcal{L}_{den}^{i}$ and $\mathcal{L}_{sup}^{i}$ are implemented as BCE losses. The ultimate loss function is formulated as their summation:
@@ 4.3.1 Effectiveness of Token-supervised Prediction Loss
Summary DUTS[101] ECSSD[102] HKU-IS[103] PASCAL-S[104] DUT-O[106] SOD[107] Metric MACs Params $S_{m}$ maxF $E_{\xi}^{\text{max}}$ MAE $S_{m}$ maxF $E_{\xi}^{\text{max}}$ MAE $S_{m}$ maxF $E_{\xi}^{\text{max}}$ MAE $S_{m}$ maxF $E_{\xi}^{\text{max}}$ MAE $S_{m}$ maxF $E_{\xi}^{\text{max}}$ MAE $S_{m}$ maxF $E_{\xi}^{\text{max}}$ MAE CNN-based PiCANet[21] 54.05 47.22 0.863 0.840 0.915 0.040 0.916 0.929 0.953 0.035 0.905 0.913 0.951 0.031 0.846 0.824 0.882 0.071 0.826 0.767 0.865 0.054 0.813 0.824 0.871 0.073 AFNet[120] 21.66 35.95 0.867 0.838 0.910 0.045 0.914 0.924 0.947 0.042 0.905 0.910 0.949 0.036 0.849 0.824 0.877 0.076 0.826 0.759 0.861 0.057 0.811 0.819 0.867 0.085 TSPOANet[121] - - 0.860 0.828 0.907 0.049 0.907 0.919 0.942 0.047 0.902 0.909 0.950 0.039 0.841 0.817 0.871 0.082 0.818 0.750 0.858 0.062 0.802 0.809 0.852 0.904 EGNet-R[18] 157.21 111.64 0.887 0.866 0.926 0.039 0.925 0.936 0.955 0.037 0.918 0.923 0.956 0.031 0.852 0.825 0.874 0.080 0.841 0.778 0.878 0.053 0.824 0.831 0.875 0.080 ITSD-R[48] 15.96 26.47 0.885 0.867 0.929 0.041 0.925 0.939 0.959 0.035 0.917 0.926 0.960 0.031 0.861 0.839 0.889 0.771 0.840 0.792 0.880 0.061 0.835 0.849 0.889 0.075 MINet-R[13] 87.11 162.38 0.884 0.864 0.926 0.037 0.925 0.938 0.957 0.034 0.919 0.926 0.960 0.029 0.856 0.831 0.883 0.071 0.883 0.769 0.869 0.056 0.830 0.835 0.878 0.074 LDF-R[19] 15.51 25.15 0.892 0.877 0.930 0.034 0.925 0.938 0.954 0.034 0.920 0.929 0.958 0.028 0.861 0.839 0.888 0.067 0.839 0.782 0.870 0.052 0.831 0.841 0.878 0.071 CSF-R2[117] 18.96 36.53 0.890 0.869 0.929 0.037 0.931 0.942 0.960 0.033 - - - - 0.863 0.839 0.885 0.073 0.838 0.775 0.869 0.055 0.826 0.832 0.883 0.079 GateNet-R[54] 162.22 128.63 0.891 0.874 0.932 0.038 0.924 0.935 0.955 0.038 0.921 0.926 0.959 0.031 0.863 0.836 0.886 0.071 0.840 0.782 0.878 0.055 0.827 0.835 0.877 0.079 MENet[122] 94.62 27.83 0.905 0.895 0.943 0.028 0.927 0.938 0.956 0.031 0.927 0.939 0.965 0.023 0.871 0.848 0.892 0.062 0.850 0.792 0.879 0.045 0.841 0.847 0.884 0.065 Generic Dense Prediction Model Mask2Former-T [96] 248.16 47.40 0.880 0.871 0.925 0.041 0.922 0.942 0.961 0.037 0.916 0.937 0.967 0.032 0.866 0.845 0.890 0.069 0.823 0.763 0.858 0.062 0.815 0.847 0.882 0.081 Transformer-based VST[45] 41.36 44.48 0.896 0.877 0.939 0.037 0.932 0.944 0.964 0.034 0.928 0.937 0.968 0.030 0.873 0.850 0.900 0.067 0.850 0.800 0.888 0.058 0.854 0.866 0.902 0.065 EBMGSOD[88] 53.38 118.96 0.909 0.900 0.949 0.029 0.941 0.954 0.972 0.024 0.930 0.943 0.971 0.023 0.877 0.856 0.899 0.061 0.858 0.817 0.900 0.051 0.866 0.881 0.915 0.056 ICON-B[87] 52.59 94.30 0.917 0.911 0.960 0.024 0.941 0.954 0.971 0.024 0.936 0.947 0.974 0.022 0.885 0.860 0.903 0.058 0.869 0.830 0.906 0.043 0.862 0.878 0.909 0.059 Ours VST-t++ 30.07 46.05 0.897 0.881 0.939 0.035 0.933 0.946 0.966 0.030 0.928 0.938 0.969 0.026 0.869 0.849 0.899 0.066 0.854 0.812 0.894 0.056 0.855 0.870 0.905 0.062 VST-T++ 28.44 53.60 0.901 0.887 0.943 0.033 0.937 0.949 0.968 0.029 0.930 0.939 0.968 0.026 0.878 0.855 0.901 0.063 0.853 0.804 0.892 0.053 0.853 0.866 0.899 0.065 VST-S++ 32.73 74.90 0.909 0.897 0.947 0.029 0.939 0.951 0.969 0.027 0.932 0.941 0.969 0.025 0.880 0.859 0.901 0.062 0.859 0.813 0.890 0.050 0.859 0.866 0.898 0.059 VST-B++ 69.30 112.23 0.923 0.915 0.957 0.024 0.950 0.962 0.977 0.021 0.941 0.951 0.976 0.021 0.889 0.867 0.907 0.056 0.871 0.827 0.900 0.044 0.872 0.886 0.910 0.056
@@ 4.4 Comparison with State-of-the-Art Methods
For RGB SOD, we compare our VST++ with 10 state-of-the-art CNN-based RGB SOD methods including PiCANet [21], AFNet [120], TSPOANet [121], EGNet-R [18], ITSD-R [48], MINet-R [13], LDF-R [19], CSF-R2 [117], GateNet-R [54], and MENet[122], as well as three state-of-the-art transformer-based RGB SOD methods: EBMGSOD [88], ICON [87], and our previous VST [45]. Moreover, we also consider one generic dense predication model, i.e. Mask2Former-T [96]. Table III shows the comparison results.
For RGB-D SOD, we adopt 15 state-of-the-art CNN-based RGB-D SOD methods, i.e., $S^{2}$ [43], PGAR [61], DANet [123], cmMS [124], ATSA [25], CMW [24], Cas-Gnn [15], HDFNet [30], CoNet [20], BBS-Net [14], JL-DCF-R [28], SPNet [29], CMINet [125], DCF [126], and SPSN [127], as well as four state-of-the-art transformer-based RGB-D SOD models, namely SwinNet [89], HRTransNet [90], EBMGSOD [88], and our previous VST [45], for comparison. Table IV and Table V report the comparison results.
@@
We additionally evaluate our VST-T++ against the retrained Mask2Former-T [96], utilizing the same backbone. The findings reveal that VST++ outperforms Mask2Former-T, highlighting the strengths of our distinctive design.